# LLM Task Interference: An Initial Study on the Impact of Task-Switch in Conversational History

### v1

@@ Appendix C Conversation History Length Ablation
This section presents an ablation study on the performance change after a task-switch for varying conversation history lengths. For each dataset in Table 4 we select four datasets (including itself), from which we use the training set to use as conversation history. The details of the prompt structure are presented in Appendix D.
C.1 Performance change with Task-switch
We compare the percentage change in metrics relative to zero-shot performance ($\mathbf{h}=\text{Ø}$, i.e. no conversation history) as a function of conversation history length $L$ and for different LLMs. Results are plot in Figures 3, 4, 5, 6, 7 for MMLU HA, MMLU AA, RT, Gigaword and TweetQA respectively. Intuitively, we would expect that when there is not a task switch, the performance would increase (assuming the training examples are aligned well with the test set). As per our discussion in Section 4.2, we observe that different models degrade on different task-switches and this is not limited by the model size.
@@ C.2 Format Failure Rate
Typically, classification tasks (MMLU HA, MMLU AA, RT) are evaluated using logits, however we chose to use a generative approach for consistency: we are evaluating the model in a conversational setting, and we do not have access to the logits exactly. Thus, we must post-process the model output to determine the class. In this, we try to give the LLM the benefit of the doubt and do our best to extract the class. For example, although the prompt requests the model to output within answer tags like "<Answer> positive </Answer>", we also accept "positive", however we do not accept "positive/negative". Due to the imperfect nature of this setup, either we may not detect the correct format, or the model generates erroneous text.
@@ Appendix E Task-Switch Sensitivity Metrics
The above are sensitivity per example, which we can use to estimate the task-switch sensitivity $\tau_{i}=\mathbb{E}[\log\ \rho_{i}]$ as per Equation 3, where the expectation is calculated over the examples and histories (for a given length $L$). We evaluate these metrics on the target task RT (rotten tomatoes) as shown in Figure 11. Figure 10(a) shows that the zero-shot sensitivity metric trends upwards for both models. This is expected for a model which does not handle task-switch well as the probability of the output with an increased conversation length decreases in comparison to the zero-shot probability. For the confidence sensitivity in Figure 10(b), we observe that Mistral-7B behaves as we expect, whereas Llama-7B becomes less confident in its output compared to having no conversation history. For the loss sensitivity metric in Figure 10(c), we observe that Llama behaves as we expect as the sensitivity remains relatively flat: as the conversation history increases, there is no significant change in the probability of outputting the reference. However, for Mistral-7b, the probability falls immediately and plateaus showing that the model was giving a very low probability mass to the reference with no conversation history. Intuitively, it is clear that both models agree in their trends only for the zero-shot sensitivity $\tau_{1}$ in Figure 10(a), hence in the main paper, we report zero-shot sensitivity as the task-switch sensitivity.

### v2

@@ Appendix C Conversation History Length Ablation
This section presents an ablation study on the performance change after a task-switch for varying conversation history lengths. For each dataset in Table 4 we select four datasets (including itself), from which we use the training set as conversation history. The details of the prompt structure are presented in Appendix D.
C.1 Task-switch Performance Change
We compare the percentage change in metrics relative to zero-shot performance ($\mathbf{h}=\text{Ø}$, i.e. no conversation history) as a function of conversation history length $L$ and for different LLMs. Results are plot in Figures 3, 4, 5, 6, 7 for MMLU HA, MMLU AA, RT, Gigaword and TweetQA respectively. When there is not a task switch, we would expect a performance increase (assuming the training examples are representative of the test set). As per our discussion in Section 4.2, we observe that different models degrade on different task-switches and this is not limited by the model size.
@@ C.2 Format Failure Rate
Typically, classification tasks (MMLU HA, MMLU AA, RT) are evaluated using logits, however we use a generative approach for consistency: we are evaluating the model in a conversational setting, and we do not have access to the logits exactly. Thus, we must post-process the model output to determine the class. In this, we try to give the LLM the benefit of the doubt and do our best to extract the class. For example, although the prompt requests the model to output within answer tags like "<Answer> positive </Answer>", we also accept "positive", but we do not accept "positive/negative". Due to the imperfect nature of this setup, either we may not detect the correct format, or the model generates erroneous text.
@@ Appendix E Task-Switch Sensitivity Metrics
The above are sensitivity per example, which we can use to estimate the task-switch sensitivity $\tau_{i}=\mathbb{E}[\log\ \rho_{i}]$ as per Equation 3, where the expectation is calculated over the examples and histories (for a given length $L$). We evaluate these metrics on the target task RT (rotten tomatoes) as shown in Figure 11. Figure 11(a) shows that the zero-shot sensitivity metric trends upwards for both models. This is expected for a model which does not handle task-switch well as the probability of the output with an increased conversation length decreases in comparison to the zero-shot probability. For the confidence sensitivity in Figure 11(b), we observe that Mistral-7B behaves as we expect, whereas Llama-7B becomes less confident in its output compared to having no conversation history. For the loss sensitivity metric in Figure 11(c), we observe that Llama behaves as we expect as the sensitivity remains relatively flat: as the conversation history increases, there is no significant change in the probability of outputting the reference. However, for Mistral-7B, the probability falls immediately and plateaus showing that the model was giving a very low probability mass to the reference with no conversation history. Intuitively, it is clear that both models agree in their trends only for the zero-shot sensitivity $\tau_{1}$ in Figure 11(a), hence in the main paper, we report zero-shot sensitivity as the task-switch sensitivity. Note finally that a further major advantage of the zero-shot sensitivity is that it is reference-free and so can be used in practice to predict the vulnerability of deployed models to different task-swithches when there is no access to the correct, reference response.
Appendix F Correlations Models, Datasets and Performance
We rank model performance against various metrics to see if there is any correlation that may help explain model performance more generally.
F.1 Task Conversation History Length
CH Task Length Llama-7B Mistral-7B
Gigaword 75 -21.35 -15.94
TweetQA 93 -15.10 -4.35
RT 108 -13.02 10.87
MMLU AA 143 -1.79 37.68
Table 11: Target Task: MMLU AA. Average length of response (length) in Conversation History (CH) compared to the performance degradation of models in the task-switch.
CH Task Length Llama-7B Mistral-7B
Gigaword 76 1.98 -0.72
TweetQA 93 2.70 -1.28
RT 108 2.38 2.83
MMLU AA 143 -5.42 1.19
Table 12: Target Task: RT. Average length of response (length) in Conversation History (CH) compared to the performance degradation of models in the task-switch.
In Table 11 and Table 12 we compare the model performance against the mean conversation history task, $T_{h}$ length, which is the average number of tokens per response in a turn in the conversation history. The model performance is taken for three different seeds with conversation history lengths $L\in\{3,6\}$. There is no observed correlation between the average length of responses for a task in a conversation and the performance degradation observed for different models.
F.2 Task Distance
In this section we aim to assess the hypothesis that the ‘distance’ between tasks can explain the extent of performance degradation in different task-switches, from the conversation history task, $T_{h}$ to the target task, $T_{t}$. Measuring distance between tasks is a multi-faceted and complex metric. Given the lack of formal task distance measures, we instead use a consensus ranking approach, where multiple powerful Large Language Models (LLMs) are required to rank the different tasks on how similar they are. For the target task RT, we queried four of the largest and most powerful models to rank the closest tasks, based on the description of each task. We consider the following LLMs: ChatGPT; Gemini Ultra (Team et al., 2024), Claude 3 Sonnet from Anthropic; and Perplexity AI. The rankings by the LLMs are given in Table 13 relative to RT. We then select an overall ranking with the greatest consensus - in this case three of the four LLMs agree perfectly in the ranking. This gives a consensus vote of ranks (relative to RT): RT (1); MMLU AA (3); TweetQA (2); and Gigaword (3). The equivalent ranks are given in Table 14 with MMLU AA as the reference task. In this case, three of the four models perfectly agree in their rankings.
Dataset ChatGPT Gemini Claude Perplexity
RT 1 1 1 1
MMLU AA 4 4 4 4
TweetQA 2 3 2 2
Gigaword 3 2 3 3
Table 13: Rank given by LLM for different datasets on how similar they are to the target task RT.
Dataset ChatGPT Gemini Claude Perplexity
RT 4 4 4 4
MMLU AA 1 1 1 1
TweetQA 2 3 2 2
Gigaword 3 2 3 3
Table 14: Rank given by LLM for different datasets on how similar they are to the target task MMLU AA.
The following tables compare the rank of the dataset distance against the mean model performance. The model performance is the %-percentage accuracy change relative to zero-shot, and the mean is taken over three seeds and over conversation history lengths $L\in\{3,6\}$.
CH-Task Rank Llama-7B Mistral-7B
RT 1 2.38 2.83
TweetQA 2 2.70 -1.28
Gigaword 3 1.98 -0.72
MMLU AA 4 -5.42 1.19
Table 15: Target Task, $T_{t}$: RT. Performance degradation (with different conversation history tasks) compared to the task rank, measuring similarity to $T_{t}$.
CH-Task Rank Llama-7B Mistral-7B
MMLU AA 1 -1.79 37.68
TweetQA 2 -15.10 -4.35
Gigaword 3 -21.35 -15.94
RT 4 -13.02 10.87
Table 16: Target Task: MMLU AA. Performance degradation (with different conversation history tasks) compared to the task rank, measuring similarity to $T_{t}$.
Overall, there appears to be only a weak correlation in some settings between the task distance and the performance degradation. This suggests that performance degradation is not only a function of the task distance, but is also an attribute of the specific model. Further analysis would be required to understand the aspects of specific models for certain task-switches that influence the level of performance degradation.
Appendix G Performance Confusion Matrix
In this section, we summarize the performance change for every pairing of task-switches from conversation history task ($T_{h}$) to target task ($T_{t}$). We present the results here for a conversation length of $L=6$ for each model separately. Tables 17, 18, 19, 20 report the results for models GPT-3.5, GPT-4, Llama-7B, Mistral-7B respectively. Each row is the performance change in the Target Task $T_{t}$. Please note that the metric for the tasks are: accuracy for MMLU AA, RT, MMLU HA, METEOR for TweetQA, and RougeL for Gigaword.
Conversation History Task
Target Task AA RT TQ GW HA
MMLU AA 19.35 6.45 6.45 -3.13 *
RT -0.22 3.00 -0.33 0.11 *
Tweet QA -13.78 -3.55 24.81 -5.69 *
Gigaword -12.10 -6.59 -3.48 67.85 *
MMLU HA * 4.73 -12.84 -8.11 20.41
Table 17: Model: GPT-3.5. Percentage % change in model performance.
Conversation History Task
Target Task AA RT TQ GW HA
MMLU AA 8.62 -13.11 -3.39 0.00 *
RT 0.76 1.74 -0.98 -0.98 *
Tweet QA 3.69 25.58 35.80 5.06 *
Gigaword 12.52 * 14.18 59.07 *
MMLU HA * 0.53 1.59 2.14 5.85
Table 18: Model: GPT-4. Percentage % change in model performance.
Conversation History Task
Target Task AA RT TQ GW HA
MMLU AA 3.57 -15.63 0.00 -10.71 *
RT -5.69 1.82 3.76 1.82 *
Tweet QA 1.68 13.37 50.74 10.17 *
Gigaword 11.76 * 11.73 158.79 *
MMLU HA * 13.86 -3.96 -0.99 25.74
Table 19: Model: Llama-7B. Percentage % change in model performance.
Conversation History Task
Target Task AA RT TQ GW HA
MMLU AA 28.57 18.18 -4.76 -19.05 *
RT 0.97 3.79 -0.87 -1.30 *
Tweet QA -0.78 7.62 30.56 9.26 *
Gigaword -3.81 2.61 3.44 78.71 *
MMLU HA * 1.63 0.81 -11.38 12.20
Table 20: Model: Mistral-7B. Percentage % change in model performance.
Appendix H Free-run Performance
In this section, we investigate the performance change for task-switch without teacher-forcing the response, $r_{k}$, in the conversation history (mentioned in our methods at the end of Section 3). We call this “free-running”. The results of the free-running experiments are shown in Fig 12.
Similar to the teacher-forced experiments, we observe that the “free-running” experiments show that model performance changes in the presence of a task-switch. We note that the results are indeed not identical because these models are not perfect at responding to the history tasks, $T_{h}$. In Fig 12, we observe that the performance can degrade based on the combination of model and task-switch. Interestingly, some task-switches allow the model to perform better than in-context examples, evident in Llama-7B for Rotten Tomatoes as $T_{t}$ and Tweet QA as $T_{h}$.
Figure 12: Percentage % change in accuracy without teacher-forcing i.e. free-running, relative to zero-shot performance (no conversation history) for increasing conversation history length $L$ and various models. The target tasks are (MMLU Abstract Algebra, Rotten Tomatoes and MMLU Human Aging).
Appendix I Random Conversation History
In this section, we study the impact on the task-switch performance with a random conversation as the history, as opposed to a specific single conversation-history task, $T_{h}$. This evaluation offers a baseline measure of a model’s general sensitivity to a change from the conversation history. This allows us to assess the extent to which the observed performance degradation from a task-switch is due to the explicit change in task, as opposed to the presence of an unrelated conversation history.
To perform this experiment we generate $K=20$ random conversation pairs $(u_{k},r_{k})$ as follows:
  • Model generates the random user utterance, $u_{k}$, starting from the <BOS> (beginning of sequence) token.
  • Model generates the system response, $r_{k}$, in a conversational setting.
This gives us a set of user-system conversation pairs, which we can use to create a ‘random’ conversation history of length $L$ by sampling pairs from this dataset.
These experiments are evaluated on the models: Llama-7B and Mistral-7B, for the target tasks ($T_{t}$): MMLU Abstract Algebra (MMLU AA) and Rotten Tomatoes (RT). The results are shown in Fig 13. We observe that both models have no task-switch performance change with $T_{t}=$ RT. However, both models have some performance change with MMLU AA as $T_{t}$; noticeably, Llama-7B suffers the most. We hypothesize that these results can be explained by the task-switch sensitivity metric $\tau$ reported in Tables 23. Table 3 shows a comparatively lower task-switch sensitivity compared to MMLU AA as $T_{t}$ in Table 2. Moreover, with MMLU AA as $T_{t}$ (Table 2), Llama-7B consistently has a higher task-switch sensitivity than Mistral-7B, which could explain the large performance change observed in Fig 13. Overall, these results confirm that observed performance degradation due to task-switches is primarily due to the change in task and not due to the presence of random/unrelated noise in the conversation history.
Figure 13: Percentage % change in accuracy with a random conversation history for increasing conversation history length $L$. Please note that in this Figure, the plotted lines correspond to the target task, not the conversation history task.